Hi, the problem i am having is with subsitution in a qudratic equation.
x^2-xy=10 (1)
x+y=8 (2)
Thank you. I worked it out using the qudratic formula but i would like to know if it's right before handing my assignment in to my teacher I need it by tommrrow morning please help.

Found 2 solutions by scott8148, mananth:
Answer by scott8148(6628)

y = 8 - x

substituting ___ x^2 - 8x + x^2 = 10 ___ 2x^2 - 8x - 10=0 ___ x^2 - 4x - 5 = 0

factoring ___ (x - 5)(x + 1) = 0

substitute to find y
___ there are two (x,y) solutions

Answer by mananth(16946)

x^2-xy=10 (1)
x+y=8 (2)
x=8-y
plug value of x in (1)
(8-y)^2-(8-y)y-10=0
64-16y+y^2-8y+y^2-10=0
2y^2-24y+54=0
/2
y^2-12y+27=0
y^2-9y-3y+27=0
y(y-9)-3(y-9)=0
(y-3)(y-9)=0
y = 3 , y=9
y=3, x=5
y=9,x=-1
...
m.ananth@hotmail.ca